OpenStudy (anonymous):

I don't know how to find the domain and range for y=4x-3 and identify any if the equation define functions

OpenStudy (mimi_x3):

D: all real values of x

OpenStudy (saifoo.khan):

wooooow. meme to the rescue.

OpenStudy (mimi_x3):

LOL

OpenStudy (anonymous):

What u mean by all real values of x how did u know that?

jimthompson5910 (jim_thompson5910):

The domain and range of any line (that is not vertical or horizontal) is the set of all real numbers.

OpenStudy (mimi_x3):

Well, for straight lines domain is all real values of x

OpenStudy (anonymous):

Then how about if the question will be y=-2(x-3)sequerd+4

OpenStudy (mimi_x3):

what's sequerd ?

OpenStudy (anonymous):

sorry miss spelled I mean squared

jimthompson5910 (jim_thompson5910):

\[\textrm{Domain is the set of all real numbers}\] \[\textrm{Range is }(-\infty,4]\]

OpenStudy (anonymous):

But I want to know how to do it not to just know the answer

OpenStudy (anonymous):

How come it's x^4 when it's (x-3)^2

OpenStudy (anonymous):

I don't know how to do it!!!! I'm waiting for someone to reply

jimthompson5910 (jim_thompson5910):

The domain of any quadratic is the set of all real numbers.

jimthompson5910 (jim_thompson5910):

The range of the quadratic \[\large y=a(x-h)^2+k\] is \[\large (-\infty, k]\] if \[\large a <0\] -------------------------- The range of the quadratic \[\large y=a(x-h)^2+k\] is \[\large [k,\infty)\] if \[\large a >0\]

jimthompson5910 (jim_thompson5910):

In this case, a=-2 and k=4, so you pick the first option described above.

OpenStudy (anonymous):

Oh, ok thank u for the help. now I got it